# Improving Retrieval Augmented Language Model with Self-Reasoning

### v2

@@ Abstract
The Retrieval-Augmented Language Model (RALM) has shown remarkable performance on knowledge-intensive tasks by incorporating external knowledge during inference, which mitigates the factual hallucinations inherited in large language models (LLMs). Despite these advancements, challenges persist in the implementation of RALMs, particularly concerning their reliability and traceability. To be specific, the irrelevant document retrieval may result in unhelpful response generation or even deteriorate the performance of LLMs, while the lack of proper citations in generated outputs complicates efforts to verify the trustworthiness of the models. To this end, we propose a novel self-reasoning framework aimed at improving the reliability and traceability of RALMs, whose core idea is to leverage reasoning trajectories generated by the LLM itself. The framework involves constructing self-reason trajectories with three processes: a relevance-aware process, an evidence-aware selective process, and a trajectory analysis process. We have evaluated our framework across four public datasets (two short-form QA datasets, one long-form QA dataset, and one fact verification dataset) to demonstrate the superiority of our method, which can outperform existing state-of-art models and can achieve comparable performance with GPT-4, while only using 2,000 training samples.
Improving Retrieval Augmented Language Model with Self-Reasoning
Yuan Xia, Jingbo Zhou, Zhenhui Shi, Jun Chen, Haifeng Huang Baidu Inc., China. {xiayuan,zhoujingbo,shizhenhui,chenjun22,huanghaifeng}@baidu.com
1 Introduction
Figure 1: An example of how self-reasoning framework generates reasoning trajectories with a relevant-aware process, an evidence-aware selective process, and a trajectory analysis process.
@@ 1 Introduction
The Retrieval-Augmented Language Model (RALM) has become a critical enhancement for Large Language Models (LLMs) by integrating external knowledge during inference. Despite their advanced capabilities in language understanding and generation Brown et al. (2020); Touvron et al. (2023), LLMs are prone to producing hallucinated and inaccurate content, particularly in knowledge-intensive tasks Ji et al. (2023). Augmenting LLMs with relevant information obtained from external sources like Wikipedia and search engines has proven effective in reducing these inaccuracies Guu et al. (2020); Lewis et al. (2020); Borgeaud et al. (2022); Izacard et al. (2022); Asai et al. (2024). This method has proven effective in mitigating the factual hallucinations inherent in LLMs Kwiatkowski et al. (2019); Petroni et al. (2021); Ram et al. (2023).
@@
Nevertheless, there are still limitations associated with RALMs, particularly in terms of reliability and traceability. First, the reliability of the retrieved information remains a significant concern. Previous studies have shown that noisy retrieval can adversely affect the performance of an LLM Menick et al. (2022); Li et al. (2023), as irrelevant data can lead to misguided responses and disturb the model’s ability to leverage its intrinsic knowledge effectively. Second, the interpretability and traceability of outputs generated by RALMs need to be improved. While RALMs incorporate retrieved documents during both training and inference phases, they may fail to cite these documents explicitly,
thus complicating the process of tracing and verifying the claims made by LLMs. To improve the retrieval robustness, recent studies have explored incorporating natural language inference (NLI) models Honovich et al. (2022) and document summarization models during inference Yoran et al. (2023); Xu et al. (2023a). However, the effectiveness of these external NLI and summarization models largely affects the overall performance of RALMs. The training and optimization of these auxiliary models require additional costs. Consequently, determining the most appropriate training and selection methods for these NLI and summarization models presents a further challenge in the application of such approaches.
@@
To address the above limitations, we propose a novel end-to-end self-reasoning framework to improve the performance of RALMs. Our intuition is that the explicit self-reasoning trajectory crafted by LLMs can improve the retrieval robustness and accuracy of question answering. During the pre-training phase, an LLM primarily focuses on knowledge acquisition, yet it does not learn to reason from retrieved documents to generate answers. To address this, a feasible approach is to incorporate reasoning trajectories into a post-training phase. Such an approach could potentially teach the model to reason and distinguish relevant and irrelevant documents, improving its response accuracy to queries. An example of how our self-reasoning framework generates reasoning trajectories is illustrated in Figure 1. In contrast, as shown in the middle part of Figure 2, the conventional RALM methods gather all documents in a non-selective manner, leading to the distraction of the LLM by irrelevant content and consequently resulting in the generation of erroneous answers.
@@
Our framework comprises three self-reasoning processes: 1) a Relevance-Aware Process (RAP), which instructs the LLM to judge the relevance between the retrieved documents and the question, 2) an Evidence-Aware Selective Process (EAP), which directs the LLM to choose and cite relevant documents, and then automatically select snippets of key sentences as evidence from the cited documents, 3) a Trajectory Analysis Process (TAP), which requires the LLM to generate a concise analysis based on all gathered self-reasoning trajectories generated by previous two processes and subsequently provide the final inferred answer. Furthermore, we propose a gradual training method by employing stage-wise masking strategies to enhance the performance of our framework.
We summarize our contributions as follows:
@@
  • We propose a novel end-to-end self-reasoning framework that improves the robustness of RALMs by leveraging reasoning trajectories generated by the LLM itself, without the need for external models or tools.
@@
  • We evaluate our framework on four public datasets (two short-form QA, one long-form QA, and one fact verification), demonstrating that our method surpasses existing state-of-art models in performance, achieving this with only using 2,000 training samples.
@@
Figure 2: An illustration of the self-reasoning framework for improving the RALMs. The upper is the basic LLMs which answer the question by inherent knowledge. The middle is the standard retrieval augmented LMs, which use retrieved documents to help answer the question. The bottom is our self-reasoning framework which uses self-generated reason trajectories to output answers.
2 Related Work
2.1 Retrieval-augmented LMs
@@
Many studies have investigated augmenting the performance of LLMs with externally retrieved information Izacard et al. (2022); Guu et al. (2020); Borgeaud et al. (2022) and some of them pre-train language models with retrieved passages. For works focusing on RALMs with citations, Menick et al. (2022); Nakano et al. (2021) instruct or train an LLM to answer questions with retrieved documents while providing citations. Gao et al. (2023b) proposes an end-to-end system to retrieve supporting evidence and generate answers with citations, while only focusing on prompting without updating their model weights. Other works instruct or fine-tune LLMs to use external tools to retrieve dynamically Schick et al. (2023); Yao et al. (2023); Jiang et al. (2023), which offers an adaptive method of when and what to search. Gao et al. (2023a) improves the attribution and factuality of language models by taking outputs of LLMs and applying a post-process retrieve-and-edit approach.
2.2 Robustness for RALMs
To improve the robustness of RALMs, Yoran et al. (2023) use a natural language inference model to filter out irrelevant documents before RALMs, and Xu et al. (2023a) use a memorization model to filter out or compress retrieved documents before using them to prompt an LLM. The concurrent work Yu et al. (2023) generates summaries of documents as notes before RALMs, and Baek et al. (2023) uses a separate small language model as a verifier to detect and correct errors in LLMs during retrieval for generating factually correct outputs. Different from the above works, ours identifies key sentences and cites relevant documents through an end-to-end framework that eliminates irrelevant ones, without dependence on external inference models.
The most similar work to ours is presented by Asai et al. (2024), who developed a method that teaches models to retrieve information using designed reflection tokens. However, this approach needs to train extra critic models and generator models to predict the reflection tokens, which requires tens of thousands of extra training samples. In contrast, our method does not rely on special tokens. Instead, we construct reasoning trajectories and then use them to directly enhance the performance of LLMs, offering a more efficient and scalable solution. More related works on LLMs for reasoning are discussed in the Appendix §$\S$§A.1.
3 Preliminary
@@
We formally define the problem of retrieval augmented generation with self-reasoning. Given a query $q$ and a corpus of documents $\mathcal{D}$, an LLM-generated answer with $m$ statements and $n$ tokens can be defined as $y=(s_{1},s_{2},\cdots,s_{m})=(w_{1},w_{2},\cdots,w_{n})$, where $s_{i}$ is the $i$-th statement and $w_{j}$ is the $j$-th token in the generated answer. In addition, for long-form QA settings, each statement $s_{i}$ should cite a list of documents $C_{i}=\{c^{(1)}_{i},c^{(2)}_{i},...\}$, where $c^{(k)}_{i}\in\mathcal{D}$. In our work, we train an LLM (e.g. LLaMA2) to first generate reasoning trajectories $\tau$ through self-reasoning and then to generate answers ${y}^{*}$ (including long-form answers and short-form answers) on condition of $\tau$. The model output is $y={\rm{concat}}(\tau,y^{*})$, which is the concatenation of $\tau$ and $y^{*}$. Note that the generations of $\tau$ and $y^{*}$ are done in a single pass within our self-reasoning framework.
4 Method
In this section, we provide a detailed implementation of the self-reasoning process and the method to construct reasoning trajectories.
@@ 4 Method
In this work, we propose a novel framework which involves three processes: 1) a Relevance-Aware Process (Sec §$\S$4.1), 2) an Evidence-Aware Selective Process (Sec §$\S$4.2), and 3) a Trajectory Analysis Process (Sec §$\S$4.3). An illustration of our self-reasoning framework for improving the RALMs is shown in Figure 2. In Sec 4.4, we demonstrate the process of data generation and quality control, and Sec 4.5 presents the details of model training.
4.1 Relevance-Aware Process
@@
In this work, we choose the DPR Karpukhin et al. (2020) and the Contriever Izacard et al. (2021) as the default retriever $R$ to recall the top-$k$ relevant documents. When presented with a question and a set of documents, people can determine whether the question is relevant to the retrieved documents. Therefore, we first instruct the model to judge the relevance between the retrieved documents $\mathcal{D}$ and the given question $q$. We further request the model to explicitly generate reasons explaining why given documents are identified as relevant. The output should include two fields as relevant and relevant reason, as depicted in Figure 2. Noting that if all of the retrieved documents are irrelevant, the model should provide an answer based on the internal knowledge acquired during its pre-training phase. We define the self-reasoning trajectories generated by RAP as $\tau_{r}$.
4.2 Evidence-Aware Selective Process
@@
When required to answer the question, people generally will first identify the crucial sentences from the provided documents, and then cite or highlight them as key points. This process of citing the document facilitates reading comprehension and can serve as a technique for combining multiple short answers to address various aspects. While people may carry out this selective process and citation instantaneously, LLMs need to formulate the self-reasoning trajectories explicitly.
@@
4.3 Trajectory Analysis Process
Finally, we consolidate all the self-reasoning trajectories ($\tau_{r}$ and $\tau_{e}$) in the previous processes together to form a chain of reasoning snippets, thereby enhancing the overall performance of the retrieval augmentation generation. Specifically, we ask the LLM to analyze the reasoning trajectories within itself and ultimately to output a concise analysis and a short answer. We instruct the LLM to output content with two fields as analysis and answer, which is shown in Figure 2. We define the self-reasoning trajectories generated by TAP as $\tau_{a}$. In this work, the analysis output is defined as a long-form answer, and the answer output is defined as a short-short answer. In Section 5.2, we further explored the performance of long-form and short-form QA settings.
4.4 Data Generation and Quality Control
@@
For the Relevance-Aware Process data generation, as manually labeling the relevant and irrelevant documents is label-intensive, we request the GPT-4 OpenAI (2023) to generate answers as ground truth. Specifically, we instruct the GPT-4 to generate labels regarding irrelevant fields, and further to output the reasons why the given documents cannot answer the question. We concatenate the given question and the retrieved documents as positive samples. For negative samples, we randomly select a different question from the training set and retrieve the top-$k$ documents related to it. These documents are then concatenated with the initial question to form negative samples. To avoid order bias in the training data, we shuffle the order of the documents.
For the EAP and TAP data generation, manually annotating the citation of the document and writing the self-reasoning process for each question is not feasible in practice. Therefore, we follow a similar process to RAP, we first instruct the GPT-4 to generate a snippet of selected documents and subsequently output the reasoning process as trajectories. The method to construct the EAP training data is the same as RAP except that the instructions for the GPT-4 are different. The details of the instructions are shown in Appendix §$\S$§A.2.
@@
For training data generation, correct and comprehensive reasoning trajectories are very important. When training an LLM, the quality of the training samples is more important than the quantity Zhou et al. (2023). As we cannot guarantee the correctness of self-reasoning trajectories and citations by the GPT-4, we develop two efficient methods to control the quality of data generation: 1) The first method is to use the off-the-shelf tools  [^1] Tools are available at https://github.com/princeton-nlp/ALCE/tree/main in Gao et al. (2023b) to automatically verify the performance of data generation for document citations. We calculate the citation precision and recall score for each training sample and filter out scores lower than our pre-defined thresholds $\delta_{p}$ and $\delta_{r}$, for citation precision and recall, respectively. 2) Second, though the validation of self-reasoning trajectories and citations generated by GPT-4 is challenging, verifying the correctness of the final answer is straightforward. Therefore, we filter out the trajectories that lead to the incorrect answers and only keep the correct ones. We totally generate 10,000 training samples by GPT-4, after the filtering strategy by quality control, we finally keep 2,000 training samples with high quality.
4.5 Model Training
We train our self-reasoning generation model $\phi$ by our constructed corpus which is augmented with self-reasoning trajectories $\tau$ using the standard language modeling objective, maximizing likelihood:
@@ 4.5 Model Training
where $\tau=\tau_{r}\oplus\tau_{e}\oplus\tau_{a}$ are the self-reasoning trajectories, $\oplus$ is the concatenation operator, $\tau_{r},\tau_{e},\tau_{a}$ are trajectories generated by above three processes respectively. $q$ is the provided question, and $y$ is the model output, including the intermediate reason trajectories and the final answer. $\mathcal{D}_{sr}$ is the training corpus augmented with self-reasoning trajectories.
@@
Specifically, we propose a stage-wise training process while we train the LLM stage by stage. In the first stage, we mask the trajectories produced by the next two stages (EAP and TAP) and train the model with a learning rate $r_{a}$. Then in the second stage, we only mask the trajectories generated by TAP and train the model with a learning rate $r_{b}$. Finally, we concatenate the reasoning trajectories from all stages and put them into a self-reasoning LLM for end-to-end training with a learning rate $r_{c}$. Hyper-parameters for training are described in Appendix §$\S$§A.4.
5 Experiments
@@
Vicuna7B Chiang et al. (2023) 28.0 55.2 62.4 24.3 45.7 40.8
Vicuna13B Chiang et al. (2023) 35.4 56.1 60.6 27.3 51.3 50.2
@@
ReAct Yao et al. (2023) - - 64.6 - - -
RECOMP Xu et al. (2023a) 38.4 - - - - -
Self-RAG7B Asai et al. (2024) 37.2 54.9 70.2 30.0 66.9 67.8
Self-RAG13B Asai et al. (2024) 38.8 55.8 72.1 31.7 70.3 71.3
@@
5.1 Datasets and Settings
To demonstrate the effectiveness of our proposed self-reasoning framework, we conduct an extensive experimental evaluation on two short-form QA datasets (NaturalQuestion Kwiatkowski et al. (2019) and PopQA Mallen et al. (2023)), one long-form QA dataset (ASQA Stelmakh et al. (2022)), and one fact verification dataset (FEVER Thorne et al. (2018)). Detailed descriptions of the datasets can be found in Appendix §$\S$§A.3. We explore off-the-shelf retrievers. We use the DPR Karpukhin et al. (2020) and the Contriever-MS MARCO Izacard et al. (2021) to retrieve the top five documents from Wikipedia.
By default, we use the DPR as a retriever for the NQ, as the DPR has been fine-tuned on the high-quality NQ data. On the PopQA, where question and answer pairs are created based on Wikipedia in 2022, therefore, for the PopQA, we use the December 2020 preprocessed Wikipedia corpus provided by Izacard et al. (2022) and use the Contriever as a retriever. For the ASQA dataset, we use GTR Ni et al. (2022) as a retrieval that corresponds to the experimental settings in Gao et al. (2023b). More settigns can be found in Appendix §$\S$§A.4.
5.2 Evaluation Metrics
@@
For long-form QA tasks, we report the EM recall as a correctness metric, and the citation recall and the citation precision for citation quality, which are the same as the metrics in Gao et al. (2023b).
@@ 5.3 Baseline Models
We evaluate strong open-source pre-trained LLMs as baseline models. For basic LLMs, we test LLaMA2-7B, LLaMA2-13B Touvron et al. (2023) and its instruction-tuned chat version LLaMA2-Chat-7B, LLaMA2-Chat-13B.
@@
To assess the performance of the retrieval augmented LMs, we use retrievers described in Section 5.1. First, we benchmark the models using the LLaMA2 and the Vicuna Chiang et al. (2023) series models for baselines. Additionally, for a fair comparison, we also include LLaMA2-FT, where LLaMA2 is fine-tuned on all the training samples generated by GPT-4 except the self-reasoning trajectories. To establish strong baselines, we compare our method against RECOMP Xu et al. (2023a), ReAct Yao et al. (2023), and Self-RAG Asai et al. (2024), all of which are trained with extra GPT-4 generated samples or external tools. We also compare our framework with GPT-4 OpenAI (2023). We include categorical comparisons with the baseline models in the Appendix §$\S$§A.5.
Figure 3: Noise robustness experiment results on three different datasets: (a) On the left is the NQ dataset, (b) in the middle is the PopQA dataset, (c) on the right is the FEVER dataset. The Self-RAG and Vicuna are 13B parameter size models.
5.4 Main Results

### v3

@@ Abstract
The Retrieval-Augmented Language Model (RALM) has demonstrated remarkable performance on knowledge-intensive tasks by integrating external knowledge during inference, which mitigates the factual hallucinations inherited in large language models (LLMs). Despite these advancements, challenges persist in the implementation of RALMs, particularly in terms of reliability and traceability. Specifically, the irrelevant document retrieval may result in unhelpful responses or even deteriorate the performance of LLMs, while the lack of appropriate citations in outputs complicates efforts to verify the trustworthiness of the models. To this end, we propose a novel self-reasoning framework aimed at improving the reliability and traceability of RALMs, whose core idea is to leverage reasoning trajectories generated by the LLM itself. The framework involves constructing self-reasoning trajectories through three processes: a relevance-aware process, an evidence-aware selective process, and a trajectory analysis process. We evaluated our framework across four public datasets (two short-form QA datasets, one long-form QA dataset, and one fact verification dataset) to demonstrate its superiority. Our method can outperform existing state-of-the-art models and achieve performance comparable with GPT-4, using only 2,000 training samples.
Introduction
Figure 1: An example of how self-reasoning framework generates reasoning trajectories.
@@ Introduction
The Retrieval-Augmented Language Model (RALM), also known as Retrieval-Augmented Generation (RAG), has become a crucial enhancement for Large Language Models (LLMs) by integrating external knowledge during inference. Despite their advanced capabilities in language understanding and generation (Brown et al. 2020; Touvron et al. 2023), LLMs are prone to producing hallucinated and inaccurate content, especially in knowledge-intensive tasks (Ji et al. 2023). Augmenting LLMs with relevant information obtained from external sources like Wikipedia and search engines has proven effective in reducing these inaccuracies (Guu et al. 2020; Lewis et al. 2020; Borgeaud et al. 2022; Izacard et al. 2022; Asai et al. 2024). This approach has proven effective in mitigating the factual hallucinations that are inherent in LLMs (Kwiatkowski et al. 2019; Petroni et al. 2021; Ram et al. 2023).
@@
Nevertheless, there are still limitations associated with RALMs, particularly concerning reliability and traceability. Firstly, the reliability of the retrieved information remains a substantial concern. Previous studies have shown that noisy retrieval can adversely affect the performance of an LLM (Menick et al. 2022; Li et al. 2023), as irrelevant data can lead to misguided responses and disturb the model’s ability to leverage its intrinsic knowledge effectively. Secondly, the interpretability and traceability of outputs generated by RALMs need to be improved. Although RALMs incorporate retrieved documents during both the training and inference phases, they may fail to explicitly cite these documents, thus complicating the process of tracing and verifying the claims made by LLMs. To improve the retrieval robustness, recent studies have explored incorporating external tools such as natural language inference (NLI) models (Honovich et al. 2022) and document summarization models during inference (Yoran et al. 2023; Xu et al. 2024). However, the effectiveness of these external tools largely influences the overall performance of RALMs. Additionally, training and optimizing these auxiliary models require additional costs. Consequently, identifying the most appropriate training and selection methods for NLI and summarization models remains a critical challenge in leveraging these approaches.
@@
To address the above limitations, we propose a novel end-to-end self-reasoning framework to improve the performance of RALMs. For convenience, we will also refer to this framework as self-reasoning RAG and use the terms interchangeably. Our intuition is that the explicit self-reasoning trajectory crafted by LLMs can improve both the retrieval robustness and accuracy in question answering. During the pre-training phase, while an LLM primarily focuses on knowledge acquisition, it does not learn to reason from retrieved documents to generate answers. To address this, a feasible approach is to incorporate reasoning trajectories into a post-training phase. Such an approach could potentially teach the model to reason and distinguish relevant and irrelevant documents, thereby enhancing its query response accuracy. An example of how our self-reasoning framework generates reasoning trajectories is illustrated in Figure 1. In contrast, as shown in the middle part of Figure 2, the conventional RALM methods gather all documents in a non-selective manner, leading to the distraction of the LLM by irrelevant content and consequently resulting in the generation of erroneous answers.
@@
Our framework constructs self-reasoning trajectories comprising three processes: 1) a Relevance-Aware Process (RAP), which instructs the LLM to judge the relevance between the retrieved documents and the question, 2) an Evidence-Aware Selective Process (EAP), which directs the LLM to choose and cite relevant documents, and then automatically select snippets of key sentences as evidence from the cited documents, 3) a Trajectory Analysis Process (TAP), which requires the LLM to synthesize a concise analysis based on all gathered self-reasoning trajectories generated by previous two processes and subsequently provide the final inferred answer. Furthermore, we propose a gradual training method by employing stage-wise masking strategies to enhance the performance of our framework. We summarize our contributions as follows:
@@
  • We propose a novel end-to-end self-reasoning framework that improves the robustness of RALMs by leveraging reasoning trajectories generated by the LLM itself, without the need for external tools.
@@
  • We evaluate our framework on four public datasets (two short-form QA, one long-form QA, and one fact verification), demonstrating that our method surpasses existing state-of-the-art models in performance using only 2,000 training samples.
@@
Figure 2: An illustration of the self-reasoning framework. The upper is the basic LLMs which answer the question by inherent knowledge. The middle is the standard retrieval augmented LMs, which use retrieved documents to help answer the question. The bottom is our self-reasoning framework which uses self-generated reason trajectories to output answers.
Related Work
Retrieval-augmented LMs
@@
Many studies have investigated augmenting the performance of LLMs with externally retrieved information (Izacard et al. 2022; Guu et al. 2020; Borgeaud et al. 2022) and some of them pre-train language models with retrieved passages. For works focusing on RALMs with citations, Menick et al. (2022); Nakano et al. (2021) instruct or train an LLM to answer questions with retrieved documents while providing citations. Gao et al. (2023b) proposes an end-to-end system to retrieve supporting evidence and generate answers with citations, while only focusing on prompting without updating their model weights. Other works instruct or fine-tune LLMs to use external tools to retrieve dynamically (Schick et al. 2023; Yao et al. 2023; Jiang et al. 2023), which offers an adaptive method of when and what to search. Gao et al. (2023a) improves the attribution and factuality of language models by taking outputs of LLMs and applying a post-process retrieve-and-edit approach.
Robustness for RALMs
To improve the robustness of RALMs, previous works can be divided into two categories. The first category utilizes retrieved documents to enhance the Chain of Thought (CoT). For example, IRCoT (Trivedi et al. 2023) iteratively uses retrieved documents to generate CoT, which is then used to retrieve further documents in subsequent steps. ReAct (Yao et al. 2023) introduces an iterative CoT paradigm that integrates reasoning with search results. However, irrelevant retrievals may produce misguided CoT, adversely affecting LLM performance (Menick et al. 2022; Li et al. 2023).
To address the issue of irrelevant retrieval information, the second category proposes using external modules to process retrieved documents during inference. For instance, Yoran et al. (2023) utilize a natural language inference model to filter out irrelevant documents, Yan et al. (2024) employ a retrieval evaluator to classify documents based on their quality, and Xu et al. (2024) and Yu et al. (2023) apply models to filter out or compress retrieved documents. Baek et al. (2023) deploy a separate small language model as a verifier to detect and correct errors in LLMs during retrieval. A method presented by Asai et al. (2024), which appears most similar to our approach, develops a technique that instructs models to retrieve information using specifically designed reflection tokens. However, this approach needs to train extra critic models and generator models to predict the reflection tokens, which requires tens of thousands of extra training samples.
Unlike the second group of works, which rely on external tools or additional modules to eliminate irrelevant information, the self-reasoning RAG method integrates self-reasoning directly into the model’s architecture, thereby enhancing the performance of LLMs and providing a more efficient and scalable solution. Further related works on LLMs for reasoning are discussed in the Appendix.
Preliminary
@@
We formally define the problem of retrieval augmented generation with self-reasoning. Given a query $q$ and a corpus of documents $\mathcal{D}$, an LLM-generated answer with $m$ statements and $n$ tokens can be defined as $y=(s_{1},s_{2},\cdots,s_{m})=(w_{1},w_{2},\cdots,w_{n})$, where $s_{i}$ is the $i$-th statement and $w_{j}$ is the $j$-th token in the generated answer. In addition, for long-form QA settings, each statement $s_{i}$ should cite a list of documents $C_{i}=\{c^{(1)}_{i},c^{(2)}_{i},...\}$, where $c^{(k)}_{i}\in\mathcal{D}$. In our work, we train an LLM (e.g. LLaMA2) to first generate reasoning trajectories $\tau$ through self-reasoning and then to generate answers ${y}^{*}$ (short-form answers) on condition of $\tau$. The model output is $y={\rm{concat}}(\tau,y^{*})$, which is the concatenation of $\tau$ and $y^{*}$. Note that the generations of $\tau$ and $y^{*}$ are done in a single pass within the self-reasoning framework.
Method
@@
Here we provide a detailed implementation of the self-reasoning process which involves three processes: 1) a Relevance-Aware Process (RAP), 2) an Evidence-Aware Selective Process (EAP), and 3) a Trajectory Analysis Process (TAP). An illustration of our self-reasoning framework is shown in Figure 2. Additionally, we outline the process of data generation and quality control, and present the specifics of model training.
Relevance-Aware Process
@@
In this work, we choose DPR (Karpukhin et al. 2020) and Contriever (Izacard et al. 2021) as default retrievers $R$ to recall the top-$k$ relevant documents. When presented with a question and a set of documents, people can determine whether the question is relevant to the retrieved documents. Therefore, we first instruct the model to judge the relevance between the retrieved documents $\mathcal{D}$ and the given question $q$. We further request the model to explicitly generate reasons explaining why given documents are identified as relevant. The output should include two fields as relevant and relevant reason, as depicted in Figure 2. If all of the retrieved documents are irrelevant, the model should provide an answer based on the internal knowledge acquired during its pre-training phase. We define the self-reasoning trajectories generated by RAP as $\tau_{r}$.
Evidence-Aware Selective Process
@@
When answering a question, people generally first identify the crucial sentences from the provided documents and then cite or highlight them as key points. This process of citing the document facilitates reading comprehension and can serve as a technique for combining multiple short answers to address various aspects. While people may carry out this selective process and citation instantaneously, LLMs need to formulate the self-reasoning trajectories explicitly.
@@
Trajectory Analysis Process
Finally, we consolidate all the self-reasoning trajectories ($\tau_{r}$ and $\tau_{e}$) in the previous processes together to form a chain of reasoning snippets, thereby enhancing the overall performance of the retrieval augmentation generation. Specifically, we ask the LLM to analyze the reasoning trajectories within itself and ultimately to output a concise analysis and a short answer. We instruct the LLM to output content with two fields as analysis and answer, which is shown in Figure 2. We define the self-reasoning trajectories generated by TAP as $\tau_{a}$. In this work, the analysis output is defined as a long-form answer, and the answer output is defined as a short-form answer. In the experiment section, we further explored the performance of long-form and short-form QA settings.
Data Generation and Quality Control
@@
For the Relevance-Aware Process data generation, as manually labeling the relevant and irrelevant documents is label-intensive, we request GPT-4 (OpenAI 2023) to generate answers as ground truth. Specifically, we instruct GPT-4 to generate labels regarding irrelevant fields, and further to output the reasons why the given documents cannot answer the question. We concatenate the given question and the retrieved documents as positive samples. For negative samples, we randomly select a different question from the training set and retrieve the top-$k$ documents related to it. These documents are then concatenated with the initial question to form negative samples. To avoid order bias in the training data, we shuffle the order of the documents.
For the EAP and TAP data generation, manually annotating the citation and writing the self-reasoning process for each question is not feasible in practice. Therefore, we follow a similar process to RAP, we first instruct GPT-4 to generate a snippet of selected documents and subsequently output the reasoning process as trajectories. The method for constructing the EAP training data is the same as RAP except that the instructions given to GPT-4 are different. The details of the instructions are shown in the Appendix.
@@
For training data generation, correct and comprehensive reasoning trajectories are very important. When training an LLM, the quality of the training samples is more important than the quantity (Zhou et al. 2023). As we cannot guarantee the correctness of self-reasoning trajectories and citations by GPT-4, we develop two efficient methods to control the quality of data generation: 1) The first method is to use the off-the-shelf tools  [^1] Tools are available at https://github.com/princeton-nlp/ALCE/tree/main in Gao et al. (2023b) to automatically verify the performance of data generation for document citations. We calculate the citation precision and recall score for each training sample and filter out scores lower than our pre-defined thresholds $\delta_{p}$ and $\delta_{r}$, for citation precision and recall, respectively. 2) Second, though the validation of self-reasoning trajectories and citations generated by GPT-4 is challenging, verifying the correctness of the final answer is straightforward. Therefore, we filter out the trajectories that lead to the incorrect answers and only keep the correct ones. We totally generate 10,000 training samples by GPT-4, after the filtering strategy by quality control, we finally keep 2,000 training samples with high quality. More details and pseudo-codes can be found in the Appendix.
Model Training
We train the self-reasoning RAG model $\phi$ by our constructed corpus which is augmented with self-reasoning trajectories $\tau$ using the standard language modeling objective, maximizing likelihood:
@@
where $\tau=\tau_{r}\oplus\tau_{e}\oplus\tau_{a}$ are the self-reasoning trajectories, $\oplus$ is a concatenation operator, $\tau_{r},\tau_{e},\tau_{a}$ are trajectories generated by above three processes respectively. $q$ is the provided question, and $y$ is the model output, including the intermediate reason trajectories and the final answer. $\mathcal{D}_{sr}$ is the training corpus augmented with self-reasoning trajectories.
@@
Specifically, we propose a stage-wise training process while we train the LLM stage by stage. In the first stage, we mask the trajectories produced by the next two stages (EAP and TAP) and train the model with a learning rate $r_{a}$. Then in the second stage, we only mask the trajectories generated by TAP and train the model with a learning rate $r_{b}$. Finally, we concatenate the reasoning trajectories from all stages and put them into a self-reasoning LLM for end-to-end training with a learning rate $r_{c}$. Hyper-parameters for training are described in the Appendix.
Experiments
@@
Vicuna7B (Chiang et al. 2023) 28.0 55.2 62.4 24.3 45.7 40.8
Vicuna13B (Chiang et al. 2023) 35.4 56.1 60.6 27.3 51.3 50.2
@@
ReAct (Yao et al. 2023) - - 64.6 - - -
RECOMP (Xu et al. 2024) 38.4 - - - - -
Self-RAG7B (Asai et al. 2024) 37.2 54.9 70.2 30.0 66.9 67.8
Self-RAG13B (Asai et al. 2024) 38.8 55.8 72.1 31.7 70.3 71.3
@@
Datasets and Settings
To demonstrate the effectiveness of our proposed self-reasoning framework, we conduct an extensive experimental evaluation on two short-form QA datasets (NaturalQuestion (Kwiatkowski et al. 2019) and PopQA (Mallen et al. 2023)), one long-form QA dataset (ASQA (Stelmakh et al. 2022)), and one fact verification dataset (FEVER (Thorne et al. 2018)). Detailed descriptions of the datasets can be found in the Appendix. We explore off-the-shelf retrievers. We use DPR (Karpukhin et al. 2020) and Contriever-MS MARCO (Izacard et al. 2021) to retrieve the top five documents from Wikipedia.
By default, we use DPR as a retriever for the NQ, as DPR has been fine-tuned on the high-quality NQ data. On the PopQA, where question and answer pairs are created based on Wikipedia in 2022, therefore, for the PopQA, we use the December 2020 preprocessed Wikipedia corpus provided by (Izacard et al. 2022) and use Contriever as a retriever. For the ASQA dataset, we use GTR (Ni et al. 2022) as a retrieval that corresponds to the experimental settings in (Gao et al. 2023b). More settings can be found in the Appendix.
Evaluation Metrics
@@
For long-form QA tasks, we report the EM recall as a correctness metric, and the citation recall and the citation precision for citation quality, which are the same as the metrics in (Gao et al. 2023b).
@@
We evaluate strong open-source pre-trained LLMs as baseline models. For basic LLMs, we test LLaMA2-7B, LLaMA2-13B (Touvron et al. 2023) and its instruction-tuned chat version LLaMA2-Chat-7B, LLaMA2-Chat-13B.
@@
First, we benchmark the models using the LLaMA2 and the Vicuna (Chiang et al. 2023) series models for baselines. Additionally, for a fair comparison, we also include LLaMA2-FT, where LLaMA2 is fine-tuned on all the training samples generated by GPT-4 except the self-reasoning trajectories. To establish strong baselines, we compare our method against RECOMP (Xu et al. 2024), ReAct (Yao et al. 2023), and Self-RAG (Asai et al. 2024), all of which are trained with extra GPT-4 generated samples or external tools. We also compare our framework with GPT-4 (OpenAI 2023). We include categorical comparisons with the baseline models in the Appendix.
Figure 3: Noise robustness experiment results on three different datasets: (a) On the left is the NQ dataset, (b) in the middle is the PopQA dataset, and (c) on the right is the FEVER dataset. The Self-RAG and Vicuna are 13B parameter size models.
Main Results